How to get a solution of 240 gallons @ 46% alcohol to 50% alcohol?
Answer by stanbon(75887)

How to get a solution of 240 gallons @ 46% alcohol to 50% alcohol?
----
Ans: Add "x" gallons of pure alcohol.
----
Equation::
alcohol + alcohol = alcohol
----
0.46*240 + 1*x = 0.50(240+x)
46*240 + 100x = 50*240 + 50x
------
50x = 4*240
x = 19.2 gallons (amt. of pure alcohol to add)
-----------------
Cheers,
Stan H.
--------------